# AFD-SLU: Adaptive Feature Distillation for Spoken Language Understanding

### v1

@@ Abstract
Fig. 1: An overview of our model AFD-SLU architecture, which consists of a Teacher Model, a Student Model, and a Dynamic Adapter. The Dynamic Adapter integrates a Residual Projection Neural Network for feature alignment and a Dynamic Distillation Coefficient for task-aware adjustment of distillation loss weights.
@@ 1 Introduction
Early research in SLU primarily focused on capturing the inherent correlation between intent detection and slot filling [2, 3, 4, 5, 6]. To further enhance model performance, subsequent work has explored knowledge distillation techniques leveraging pre-trained language models [7]. Li et al. [8] proposed a multi-teacher framework that transfers knowledge from multiple pre-trained models under limited supervision. However, such methods are often tailored to specific domains or languages, exhibiting limited cross-domain adaptability and poor generalization. With the rapid advancement of Large Language Models (LLMs), recent studies have begun to harness their strong generalization and few-shot capabilities for SLU [9, 10, 11]. Notably, Yin et al. [12] reformulated slot filling as an entity recognition task and proposes a ”Chain of Intent” mechanism to support multi-intent understanding within an LLM-based framework. Despite their promising performance, LLMs entail significant memory and computational costs, which poses challenges for their deployment in real-world, resource-constrained SLU applications [13, 14].
@@
To address these limitations, we propose Adaptive Feature Distillation for Spoken Language Understanding (AFD-SLU), a lightweight framework for dynamic feature knowledge distillation. It comprises three components: a teacher model, a student model, and a dynamic adapter. The teacher model is a General Text Embeddings (GTE) model [18] derived from a high-performance LLM, which provides generalizable semantic features while avoiding the heavy computational cost of direct LLM inference. To the best of our knowledge, this is the first work to employ GTE as a teacher model for SLU. The student model adopts a joint architecture for intent detection and slot filling, while the dynamic adapter integrates a Residual Projection Neural Network (RPNN) to align feature spaces and a Dynamic Distillation Coefficient (DDC) to adaptively adjust the distillation weight. This design improves the student’s generalization while remaining highly efficient. Experiments on the ProSLU benchmark demonstrate that AFD-SLU achieves state-of-the-art performance, with 95.67% intent accuracy, 92.02% slot F1 score, and 85.50% overall accuracy. Ablation studies verify the effectiveness of each module, and further experiments show the adaptability of our framework to different teacher models.
@@ 2.3.2 Dynamic Distillation Coefficient
where $\lambda_{\text{initial}}$ and $\lambda_{\text{final}}$ denote the initial and final weights, $e$ is the current epoch, and $E$ is the total number of epochs.
This cosine annealing schedule gradually reduces the influence of the teacher model, enabling the student model to increasingly focus on task-specific learning in the later stages of training. The overall training objective is formulated as:
@@
where $\mathcal{L}_{\text{task}}$ is the task-specific loss for joint intent detection and slot filling, and $\mathcal{L}_{\text{distill}}$ is the feature-level distillation loss.
@@ 3.1 Experimental Setup
Implementation Details. We train our model on a NVIDIA RTX 4090D GPU equipped with 48 GB of memory. We optimize the model using the Adam and train it for 50 epochs, applying a dropout rate of 0.4. The hyperparameters $\lambda_{\text{initial}}$ and $\lambda_{\text{final}}$ are set to 0.1 and 0.7, respectively. We evaluate our method with both JPIS [16] and PRO-HAN [17] serving individually as student models, and employ the gte_Qwen2-1.5B-instruct [18] as teacher model. For JPIS, we use a batch size of 1 and a learning rate of $2\times 10^{-4}$; for PRO-HAN, the batch size is 64 and the learning rate is $1\times 10^{-3}$.
@@ 3.2 Main Results
We apply our AFD-SLU framework to the above student–teacher settings. As shown in Table 1, integrating AFD improves performance across all evaluation metrics. For JPIS, applying AFD yields absolute gains of 2.25% in intent accuracy, 2.75% in slot F1 score, and 3.20% in overall accuracy. Similarly, PRO-HAN-AFD achieves 95.67% intent accuracy and 92.02% slot F1 score, outperforming the original model by 3.58% and 2.01%, respectively, while maintaining comparable overall accuracy. These results demonstrate that AFD-SLU enhances model performance under low-resource conditions. The consistent gains confirm that our method effectively transfers semantic knowledge from GTE models to student models, improving both intent detection and slot filling without increasing inference overhead.
@@
Model Intent (Acc) Slot (F1) Overall (Acc) SF-ID [2] 83.24 73.70 68.36 Slot-Gated [3] 83.24 74.18 69.11 Bi-Model [4] 81.92 77.76 73.45 AGIF [5] 81.54 80.57 74.95 Stack-Propagation [6] 83.99 81.08 78.91 General-SLU [15] 85.31 83.27 79.10 GL-GIN [19] 85.69 82.70 79.28 JPIS [16] 87.95 85.76 82.30 JPIS-AFD (Ours) 90.20 88.51 85.50 PRO-HAN [17] 92.09 90.01 85.31 PRO-HAN-AFD (Ours) 95.67 92.02 85.50
@@
Model Intent (Acc) Slot (F1) Overall (Acc) JPIS-AFD 90.20 88.51 85.50 w/o RPNN 87.95 86.13 83.43 w/o DDC 87.59 85.96 83.24
@@ 3.3 Ablation study
We perform an ablation study to assess the contributions of key components in our framework. As shown in Table 2, removing the RPNN results in performance drops of 2.01% in intent accuracy, 2.38% in slot F1 score, and 2.07% in overall accuracy compared to the JPIS-AFD model. This suggests that RPNN is essential for aligning the embeddings features between student and teacher models, indicating that a simple linear transformation is insufficient for effective feature alignment. Similarly, excluding the DDC leads to reductions of 2.61%, 2.55%, and 2.26% in the same metrics. This highlights the importance of adaptively adjusting the distillation weight to balance the student’s intrinsic learning and the knowledge transferred from the teacher during training. Nevertheless, even without these components, the models still outperform the original JPIS, demonstrating the overall strength and robustness of the proposed AFD-SLU framework. These results confirm that both RPNN and DDC are essential to the effectiveness of our method, working together to significantly enhance performance on joint SLU tasks.
@@ 3.4 Analysis
We evaluated the impact of different GTE models used as teacher models, with results shown in Table 3. We can conclude that: Smaller or moderately sized teacher models that are well-aligned with the task, such as GTE-Qwen2-1.5B-instruct and Qwen3-Embedding-0.6B, are more effective in guiding the student model through knowledge distillation. These models achieve better performance across intent recognition, slot filling, and overall accuracy. In contrast, overly large models tend to overfit when trained on small datasets like ProSLU, leading to suboptimal generalization and lower performance. Additionally, general-purpose pre-trained models such as GTE-large-zh may lack sufficient task-specific understanding, which limits their effectiveness as teacher models. These findings emphasize the importance of selecting a teacher model that is not only powerful but also well-suited to the task and dataset scale.
Table 3: Impact of Different General Text embeddings LLMs as Teacher Models on Knowledge Distillation Performance. Using JPIS as the student model.

### v2

@@ Abstract
Fig. 1: An overview of our model AFD-SLU.
@@ 1 Introduction
Early research in SLU primarily focused on capturing the inherent correlation between intent detection and slot filling [2, 3, 14, 10, 8]. To further enhance model performance, subsequent work has explored knowledge distillation techniques leveraging pre-trained language models [13]. Li et al. [5] proposed a multi-teacher framework that transfers knowledge from multiple pre-trained models under limited supervision. However, task-level distillation are tailored to specific domains or languages, exhibiting limited cross-domain adaptability and generalization. With the rapid advancement of Large Language Models (LLMs), recent studies have begun to harness their strong generalization and few-shot capabilities for SLU [19, 11, 16]. Notably, Yin et al. [18] reformulated slot filling as an entity recognition task and proposes a ”Chain of Intent” mechanism to support multi-intent understanding within an LLM-based framework. Despite their promising performance, LLMs entail significant memory and computational costs, which poses challenges for their deployment in real-world, resource-constrained SLU applications [1, 4].
@@
To address these limitations, we propose Adaptive Feature Distillation for Spoken Language Understanding (AFD-SLU), a lightweight framework for dynamic feature knowledge distillation. It comprises three components: a teacher model, a student model, and a dynamic adapter. The teacher model is a General Text Embeddings (GTE) model [6] derived from a high-performance LLM, which provides generalizable semantic features while avoiding the heavy computational cost of direct LLM inference. To the best of our knowledge, this is the first work to employ GTE as a teacher model for SLU. The student model adopts a joint architecture for intent detection and slot filling, while the dynamic adapter integrates a Residual Projection Neural Network (RPNN) to align feature spaces and a Dynamic Distillation Coefficient (DDC) to adaptively adjust the distillation weight. Experiments on the ProSLU benchmark demonstrate that AFD-SLU achieves state-of-the-art performance. Ablation studies verify the effectiveness of each module, and further experiments show the adaptability of our framework to different teacher models. All codes for this work are publicly available at https://github.com/XieyRock/AFD-SLU.
@@ 2.3.2 Dynamic Distillation Coefficient
where $\lambda_{\text{initial}}$ and $\lambda_{\text{final}}$ denote the initial and final weights, $e$ is the current epoch, and $E$ is the total number of epochs. This cosine annealing schedule gradually reduces the influence of the teacher model, enabling the student model to increasingly focus on task-specific learning in the later stages of training.
$\mathcal{L}_{\text{task}}=\alpha\mathcal{L}_{\text{ID}}+(1-\alpha)\mathcal{L}_{\text{SF}}$ (7)
where $\mathcal{L}_{\text{ID}}$ and $\mathcal{L}_{\text{SF}}$ denote the intent loss and the slot loss.
The overall training objective is formulated as:
@@ 3.1 Experimental Setup
Implementation Details. We train our model on a NVIDIA RTX 4090D GPU equipped with 48 GB of memory. We optimize the model using the Adam and train it for 50 epochs, applying a dropout rate of 0.4. The dynamic parameter $\lambda$ was selected based on comparative validation. Its initial and final values were set to $\lambda_{\text{initial}}=0.1$ and $\lambda_{\text{final}}=0.7$, respectively. We evaluate our method with both JPIS [7] and PRO-HAN [12] serving individually as student models, and employ the gte_Qwen2-1.5B-instruct [6] as teacher model. For JPIS, we use a batch size of 1 and a learning rate of $2\times 10^{-4}$; for PRO-HAN, the batch size is 64 and the learning rate is $1\times 10^{-3}$.
@@ 3.2 Main Results
We apply our AFD-SLU framework to the above student–teacher settings. As shown in Table 1, integrating AFD improves performance across all evaluation metrics. For JPIS, applying AFD yields absolute gains of 2.25% in intent accuracy, 2.75% in slot F1 score, and 3.20% in overall accuracy. Similarly, PRO-HAN-AFD achieves 95.67% intent accuracy and 92.02% slot F1 score, outperforming the original model by 3.58% and 2.01%, respectively, while maintaining comparable overall accuracy. These results demonstrate that AFD-SLU significantly enhances model performance under low-resource conditions, while keeping the model parameters and the training data scale nearly unchanged. The consistent gains confirm that our method effectively transfers semantic knowledge from GTE models to student models, improving both intent detection and slot filling without increasing inference overhead.
@@
Model Intent (Acc) Slot (F1) Overall (Acc) Params SF-ID [2] 83.24 73.70 68.36 <<1M Slot-Gated [3] 83.24 74.18 69.11 <<1M Bi-Model [14] 81.92 77.76 73.45 $\sim$10M AGIF [10] 81.54 80.57 74.95 $\sim$1M Stack-Propagation [8] 83.99 81.08 78.91 $\sim$10M General-SLU [17] 85.31 83.27 79.10 $\sim$1M GL-GIN [9] 85.69 82.70 79.28 $\sim$1M JPIS [7] 87.95 85.76 82.30 $\sim$2M JPIS-AFD (Ours) 90.20 88.51 85.50 $\sim$2M PRO-HAN [12] 92.09 90.01 85.31 $\sim$40M PRO-HAN-AFD (Ours) 95.67 92.02 85.50 $\sim$40M
@@
Model Intent (Acc) Slot (F1) Overall (Acc) JPIS-AFD 90.20 88.51 85.50 RPNN(Linear) 87.95 86.13 83.43 RPNN(Deep) 88.70 87.13 83.99 w/o DDC 87.59 85.96 83.24
@@ 3.3 Ablation study
We perform an ablation study to assess the contributions of key components in our framework. As shown in Table 2, RPNN (Linear) applies a single linear projection and results in performance drops of 2.01% in intent accuracy, 2.38% in slot F1 score, and 2.07% in overall accuracy compared to the JPIS-AFD model. Meanwhile, RPNN (Deep) introduces a deeper nonlinear projection network, also leads to reduced performance.This observation suggests that effective feature alignment benefits from a carefully balanced RPNN design, whereas both overly simple linear transformations and excessively complex projection networks are suboptimal. Similarly, excluding the DDC leads to reductions of 2.61%, 2.55%, and 2.26% in the same metrics. This highlights the importance of adaptively adjusting the distillation weight to balance the student’s intrinsic learning and the knowledge transferred from the teacher during training. These results confirm that both RPNN and DDC are essential to the effectiveness of our method, working together to significantly enhance performance on joint SLU tasks.
@@ 3.4 Analysis
We evaluated the impact of different GTE models used as teacher models, with results shown in Table 3. We can conclude that: Smaller or moderately sized teacher models that are well-aligned with the task, such as GTE-Qwen2-1.5B-instruct and Qwen3-Embedding-0.6B, are more effective in guiding the student model through knowledge distillation. These models achieve better performance across intent recognition, slot filling, and overall accuracy. In contrast, overly large models tend to overfit when trained on small datasets such as ProSLU, exhibiting high initial accuracy but subsequently suffering from overfitting, which leads to suboptimal generalization and degraded performance. Additionally, pre-trained models such as GTE-large-zh may lack sufficient task-specific understanding, which limits their effectiveness as teacher models. These findings emphasize the importance of selecting a teacher model that is not only powerful but also well-suited to the task and dataset scale.
Table 3: Impact of different general text embedding models used as teacher models on knowledge distillation performance, with JPIS as the student model.